An employment agency pays heavy equipment operators $136.00 a day and general laborers $81 a day. If thirty-one people were hired and the payroll was $3391, how many heavy equipment operators were employed? How many laborers?
Thanks in advance for your helpThis question is from textbook
: I need help solving this problem,
An employment agency pays heavy equipment operators $136.00 a day and general laborers $81 a day. If thirty people were hired and the payroll was $3391, how many heavy equipment operators were employed? How many laborers?
Thanks in advance for your help

Answer by checkley77(12844)

136X+81(31-X)=3391
136X+2511-81X=3391
55X=3391-2511
55X=880
X=880/55
X=16 IS THE NUMBER OF HEAVY EQUIPMENT OPERATORS.
31-16=15 IS THE NUMBER OF LABORERS.
PROOF:
136*16+81*15=3391
2176+1215=3391
3391=3391